OpenStudy (anonymous):

If log base b of 2 =x and log base b of 3 =y, evaluate the following in terms of x and y: a). log base b of 648 b). log base b of (8/81) c). log base b of 81/log base b of 4

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

we have to write 348 in term of 2 and 3

OpenStudy (misty1212):

by the miracle of the wolf, it turns out that \[648=2^3\times 3^4\] http://www.wolframalpha.com/input/?i=648

OpenStudy (misty1212):

that means \[\log_b(648)=\log_b(2^3\times 3^4)\] now we use two properties of the log

OpenStudy (anonymous):

so would the answer be x^3 times y^4?

OpenStudy (misty1212):

oh no lets go slow

OpenStudy (anonymous):

brb

OpenStudy (misty1212):

\[\log(AB)=\log(A)+\log(B)\] so the first step it two write \[\log_b(2^3\times 3^4)=\log(2^2)+\log(3^4)\]

OpenStudy (misty1212):

ok i will wait

OpenStudy (misty1212):

typo there anyways, it should be \[\log_b(2^3\times 3^4)=\log(2^3)+\log(3^4)\]

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

hmm what the problem, didn't misty help you?

OpenStudy (anonymous):

she did but then i left so she didnt finish helping me

OpenStudy (xapproachesinfinity):

the key is the write the number as power with the base 2 and 3

OpenStudy (xapproachesinfinity):

picking from where she left you need to use the power property of logs

OpenStudy (xapproachesinfinity):

to bring down the power

OpenStudy (anonymous):

3log2 + 4log3

OpenStudy (xapproachesinfinity):

yes

OpenStudy (xapproachesinfinity):

and log2=x log3=y so it would be 3x+4y

OpenStudy (anonymous):

ok i get it now thnx

OpenStudy (xapproachesinfinity):

no problem it is the same thing with the other one write them into powers then use the same method

OpenStudy (anonymous):

ok thank you

OpenStudy (xapproachesinfinity):

i guess woflram would help if you are lazy hehe :)

OpenStudy (anonymous):

aha yass xp

OpenStudy (xapproachesinfinity):

oh the second is letter different first you use anther property of log \[\log(\frac{a}{b})=\log a- \log b\]

OpenStudy (xapproachesinfinity):

so =log8-log81 =log 2^3 -log 3^4

OpenStudy (anonymous):

yeah i was doing that

OpenStudy (xapproachesinfinity):

:)

OpenStudy (anonymous):

can you help with the last one?

OpenStudy (xapproachesinfinity):

eh that should be easy log 3^4 / log 2^2

OpenStudy (anonymous):

ok i got 4y/2x